[Brahms-l] Shift report 20050221 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Feb 21 2005 - 23:44:58 EST
Shifter leader was EBJ
Shifters were 


1. There was good beam for 4 hours at the beginning of the shift, but it took them three hours to get the machine back up after a scheduled dump.  

2.  I took data with the FS at 8deg B2442 and the MRS at 40deg A1050.  I also took a tpc and a non-tpc pedetal run.

3. The detector seemd to be working fine.

4. I switched the FS magnets to A polarity and we are taking data at 8deg A2442 right now.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 21 23:45:17 2005

This archive was generated by hypermail 2.1.8 : Mon Feb 21 2005 - 23:45:30 EST